rogers age is 1 1/3 times stevens age.eight years ago rogans age was twice stevens age.find rogers age
Found 2 solutions by ptaylor, solver91311:
Answer by ptaylor(2198)

Let x=Steven's age; x-8=Stevens age 8 years ago
Then (4/3)x=Roger's age; (4/3)x-8=Roger's age 8 years ago
Now we are told that:
(4/3)x-8=2(x-8) multiply each term by 3
4x-24=6(x-8) get rid of parens
4x-24=6x-48 add 24 to and subtract 6x from each side
4x-6x-24+24=6x-6x+24-48 collect like terms
-2x=-24 divide each side by -2
x=12-------Steven's age
(4/3)x=(4/3)*12=16 Roger's age
Ck
8 years ago, Roger was twice Steven's age:
2(12-8)=16-8
8=8
Hope this helps---ptaylor

Answer by solver91311(24713)



Let represent Roger's age. Let represent Steven's age.



I'm going to go way out on a limb and presume that "Roger" and "Rogan" are the same guy who's name was mangled by your poor typing and proofreading skills. In the meantime, please consider how discourteous it is of you not to proofread your work carefully before posting.



Solve the system for .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism